OpenStudy (anonymous):

How to solve?? x + 1 divided by 5 = 2x - 2 divided by 6

OpenStudy (kl0723):

have you tried moving both terms to the same side and setting equal to zero?

OpenStudy (anonymous):

This is what I got so far but I think I'm wrong. I would first start out by cross multiplying which would leave me with 5(2x-2) = 6(x+1), then I would distribute and get 7x - 2 = 6x + 1 .

OpenStudy (kl0723):

ok, multiply both side by 30 to get rid of the denominators and leave a basic equation, such as 6(x+1) = 5(2x-2)

OpenStudy (anonymous):

I don't understand...

OpenStudy (kl0723):

30(x+1/5) = 30(2x-2/6)... then you should get 6(x+1) = 5(2x-2) and then distribute on both sides

OpenStudy (anonymous):

I did that, basically. then I would distribute and get 6x + 1 = 7x - 2

OpenStudy (kl0723):

6(x+1) = 5(2x-2) ... Distribute to all terms and you get 6x + 6 = 10x -10

OpenStudy (kl0723):

you should multiply 6 times x = 6x and then 6 times 1 = 6... all together is 6x+ 6... then do the same on the other side

OpenStudy (anonymous):

I see where I messed up now, I didn't distribute to all terms.

OpenStudy (anonymous):

How to I solve for x now?

OpenStudy (kl0723):

Yes... I am not giving up an answer so you can understand the process.. it is important... once you have done that you should get 6x + 6 = 10x - 10

OpenStudy (kl0723):

move all the x values to the left side and that yields: 6x - 10x = - 6 - 10

OpenStudy (kl0723):

you should come up with: -4x = -16 and then isolate x by dividing it by -4 and the divide -4 on the other side too, so you keep a balanced equation

OpenStudy (kl0723):

-4/-4 = x on left side... and -16/-4 = 4 on right side, so x=4 :)

OpenStudy (kl0723):

Kelly, did you understand it?